OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (4, 2).

OpenStudy (anonymous):

y=mx+b 2=2*4+b 2=8+b b=-6 y=2x-6

OpenStudy (anonymous):

ha, thanks. your the first person who actually showed me not just told me.

OpenStudy (anonymous):

you are welcome :)